OpenStudy (anonymous):

A square pyramid has volume 147cm^3. If the height of the pyramid is 9 cm, what is the length of each side of the square base?

OpenStudy (anonymous):

1/3 X 9 X a^2 = 147 a^2 = 147/3 = 49 a = side of base = 7 cms